OpenStudy (mathmath333):

Find two natural numbers whose sum is 85 and the least common multiple is 102.

OpenStudy (gorv):

let first number is= x other wll be =85-x x*(85-x)=102 solve it further

OpenStudy (mathmath333):

it dont think it gives integers

OpenStudy (mathmath333):

http://www.wolframalpha.com/input/?i=x*%2885-x%29%3D102

OpenStudy (gorv):

factorise 102

OpenStudy (gorv):

that will be 2,3,17

OpenStudy (gorv):

102 = 2*3*17 17X + 6Y = 85 ( no solution coz 17 is a factor of 85 and 6 is not a factor 17) 34X + 3Y = 85 (no solution) 51X + 2Y = 85 (X = 1 and Y = 17) so , numbers are 51 and 34

ganeshie8 (ganeshie8):

For a general method without bruteforce, you may try something like below : Say the required two numbers are \(A\) and \(B\) We're given : \[\large A+B = 85 = 5\times 17\tag{1}\] \[LCM(A,B) = 102 = 2\times 3\times 17 \] \(\gcd(A,B) \) divides both \(A\) and \(B\), it has to divide \(\large A+B\) also; that makes \(\gcd(A,B) = 17\) using identity : \(A*B = \gcd(A,B) \times LCM(A,B) \), we have : \(\large AB =17\times 102 \tag{2}\) you can solve \((1), (2)\)